# Constraints on Pre-Big-Bang Cosmology from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

### v1

@@ 3 Data analysis
In this section, we present the methodology employed to constrain the SGWB in the PBB model using GW data from the first three observing runs of the Advanced LIGO and Virgo detectors following our previous work [52, 53]. The detector network consists of the LIGO-Hanford, LIGO-Livingston, and Virgo detectors, each labeled by the index $I={H,L,V}$. The analysis spans the frequency range of $20\sim 1726$ Hz, determined by the detector sensitivity and sampling rate. The time-series output, $s_{I}(t)$, of each detector is converted into the frequency domain using a Fourier transform, resulting in $\tilde{s}_{I}(f)$.
@@
To search for the SGWB signal, we utilize the cross-correlation statistic $\hat{C}^{IJ}(f)$ for each detector pair (baseline) $IJ$, as given by [54, 55]
$\hat{C}^{IJ}(f)=\frac{2}{T}\frac{\operatorname{Re}[\tilde{s}_{I}^{\star}(f)\tilde{s}_{J}(f)]}{\gamma_{IJ}(f)S_{0}(f)},$ (3.1)
@@
where $T$ represents the observation time, $\gamma_{IJ}(f)$ denotes the normalized overlap reduction function [55] that accounts for the geometric sensitivity of the detector pair, and $S_{0}(f)=(3H_{0}^{2})/(10\pi^{2}f^{3})$ is a normalization factor related to the critical energy density of the Universe. The overlap reduction function is normalized such that $\gamma_{IJ}(0)=1$ for co-located and co-aligned detectors.
@@
$\sigma_{IJ}^{2}(f)\approx\frac{1}{2T\Delta f}\frac{P_{I}(f)P_{J}(f)}{\gamma_{IJ}^{2}(f)S_{0}^{2}(f)},$ (3.2)
@@
$\beta$ Uniform$[-1,3]$ $0.42^{+1.84}_{-0.63}$
$\log_{10}z_{s}$ Uniform$[0,25]$ $11.3^{+8.0}_{-6.6}$
$\log_{10}z_{d}$ Uniform$[0,20]$ $6.7^{+7.8}_{-4.9}$
$\log_{10}z_{\sigma}$ Uniform$[0,18]$ $3.7^{+7.2}_{-3.3}$
@@
We perform a Bayesian analysis to search for the SGWB signal originating from the PBB model, using the publicly available, model-independent cross-correlation spectra $\hat{C}^{IJ}(f)$ data [56] from the first three observing runs of Advanced LIGO and Advanced Virgo detectors. To estimate the parameters of the SGWB model arising from the PBB model, we construct a likelihood function by combining the cross-correlation spectra from all detector pairs $IJ$ [57]:
$p(\hat{C}^{IJ}(f_{k})|\boldsymbol{\theta})\propto\exp\left[-\frac{1}{2}\sum_{IJ}\sum_{k}\left(\frac{\hat{C}^{IJ}(f_{k})-\Omega_{\mathrm{M}}(f_{k}|\boldsymbol{\theta})}{\sigma_{IJ}^{2}(f_{k})}\right)^{2}\right],$ (3.3)
where $\boldsymbol{\theta}$ represents the set of parameters characterizing the SGWB model, denoted by $\Omega_{\mathrm{M}}(f|\boldsymbol{\theta})$. The likelihood assumes that the cross-correlation spectra $\hat{C}^{IJ}(f_{k})$ follow a Gaussian distribution in the absence of a signal. The sum runs over all frequency bins $k$ and detector pairs $IJ$, with $\sigma_{IJ}^{2}(f_{k})$ being the variance of the cross-correlation statistic at each frequency bin. Using Bayes’ theorem, we express the posterior distribution of the model parameters as
$p(\boldsymbol{\theta}|C_{k}^{IJ})\propto p(C_{k}^{IJ}|\boldsymbol{\theta})\,p(\boldsymbol{\theta}),$ (3.4)
where $p(\boldsymbol{\theta})$ represents the prior distribution on the parameters.
@@
$\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}=\frac{p(\hat{C}^{IJ}|\text{Model with SGWB signal})}{p(\hat{C}^{IJ}|\text{Pure noise model})}=\frac{\int p(\hat{C}^{IJ}|\boldsymbol{\theta}_{\mathrm{GW}})\,p(\boldsymbol{\theta}_{\mathrm{GW}})\,\mathrm{d}\boldsymbol{\theta}_{\mathrm{GW}}}{\mathcal{N}}.$ (3.5)
The numerator denotes the marginal likelihood of the model incorporating the SGWB signal, which is calculated by integrating the likelihood $p(\hat{C}^{IJ}|\boldsymbol{\theta}_{\mathrm{GW}})$ multiplied by the prior $p(\boldsymbol{\theta}_{\mathrm{GW}})$ across the parameter space $\boldsymbol{\theta}_{\mathrm{GW}}$. The denominator $\mathcal{N}$ represents the evidence for the pure noise model, obtained by setting $\Omega_{\mathrm{M}}(f)=0$ in Eq. (3.3). It represents the probability of observing the data given the model that assumes only the presence of noise. The Bayes factor provides a quantitative measure of the relative support for the model with the SGWB signal compared to the pure noise model. A value of $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}>1$ indicates that the data favor the model with the SGWB signal over the pure noise model. The strength of the evidence can be interpreted using a standard scale, such as the Jeffreys scale [58], where Bayes factors $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}$ exceeding 3, 10, 30, and 100 indicate substantial, strong, very strong, and decisive evidence in favor of the model with the SGWB signal, respectively.
The free parameters in our analysis are $\boldsymbol{\theta}_{\mathrm{GW}}\equiv(\beta,z_{s},z_{d},z_{\sigma})$. We list the priors of the four free parameters in Table 1. These parameters are subject to the following theoretical constraints [43]:
$1\lesssim z_{\sigma}<z_{d}<z_{s},$ (3.6)
$\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)+\frac{3}{2}\log_{10}z_{d}-\frac{7}{2}\log_{10}z_{\sigma}\leqslant 0,$ (3.7)
$\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)-3\log_{10}z_{d}+\log_{10}z_{\sigma}>-42-\log_{10}4,$ (3.8)
$\log_{10}z_{s}<26-\log_{10}9+\frac{1}{2}\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)+\frac{1}{2}\left(\log_{10}z_{\sigma}-\log_{10}z_{d}\right).$ (3.9)
We conduct the Bayesian analysis using the Bilby package [59, 60], employing the dynamic nested sampling algorithm implemented in Dynesty [61] with $1024$ live points to ensure adequate sampling of the parameter space.
4 Results and discussions
@@
Our analysis reveals no statistically significant evidence for the presence of a PBB SGWB signal in the Advanced LIGO and Advanced Virgo data. The Bayes factor between the PBB model and the noise-only model is found to be $0.03$, indicating “very strong” preference for the noise-only model. Consequently, we establish a lower limit for the parameter $\beta$ as $\beta\gtrsim-0.19$ at 90% confidence level. The theoretically viable range for this parameter is $30\leq\beta<$ 3, where the lower bound arises from the requirement of growing string coupling necessary for a smooth bounce transition [62, 63, 64], and the upper bound prevents background instabilities [65]. Notably, our observational lower limit of $\beta\gtrsim-0.19$ is compatible with the theoretical constraint $\beta\geq 0$, indicating that while we do not detect a SGWB signal, our results remain consistent with the basic theoretical framework of PBB cosmology.
Our findings underscore the potential of GW experiments in probing alternative cosmological models and exploring the pre-big-bang era. As the sensitivity of GW detectors continues to improve, we anticipate placing even more stringent constraints on PBB models or potentially detecting the PBB SGWB signal. The upcoming observing runs of Advanced LIGO and Advanced Virgo, along with future detectors such as the Einstein Telescope [66] and Cosmic Explorer [67], will play pivotal roles in this endeavor, opening new avenues for investigating the earliest stages of the Universe and testing the fundamental principles of string theory and quantum gravity.
Acknowledgments
QT is supported by the National Natural Science Foundation of China (Grants No. 12405055 and No. 12347111), the China Postdoctoral Science Foundation (Grant No. 2023M741148), the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20240458). ZCC is supported by the National Natural Science Foundation of China under Grant No. 12405056 and the innovative research group of Hunan Province under Grant No. 2024JJ1006. YW is supported by the National Natural Science Foundation of China under Grant No. 12405057. LL is supported by the National Natural Science Foundation of China Grant under Grant No. 12433001.

### v2

@@ III Data analysis
In this section, we present the methodology employed to constrain the SGWB in the PBB model using GW data from the first three observing runs of the Advanced LIGO and Virgo detectors following our previous work Chen and Liu (2024); Wu et al. (2025). The detector network consists of the LIGO-Hanford, LIGO-Livingston, and Virgo detectors, each labeled by the index I=H,L,VI={H,L,V}. The analysis spans the frequency range of $20\sim 1726$ Hz, determined by the detector sensitivity and sampling rate. The time-series output, $s_{I}(t)$, of each detector is converted into the frequency domain using a Fourier transform, resulting in $\tilde{s}_{I}(f)$.
@@
To search for the SGWB signal, we utilize the cross-correlation statistic $\hat{C}^{IJ}(f)$ for each detector pair (baseline) $IJ$, as given by Romano and Cornish (2017); Allen and Romano (1999)
$\hat{C}^{IJ}(f)=\frac{2}{T}\frac{\operatorname{Re}[\tilde{s}_{I}^{\star}(f)\tilde{s}_{J}(f)]}{\gamma_{IJ}(f)S_{0}(f)},$ (14)
@@
where $T$ represents the observation time, $\gamma_{IJ}(f)$ denotes the normalized overlap reduction function Allen and Romano (1999) that accounts for the geometric sensitivity of the detector pair, and $S_{0}(f)=(3H_{0}^{2})/(10\pi^{2}f^{3})$ is a normalization factor related to the critical energy density of the Universe. The overlap reduction function is normalized such that $\gamma_{IJ}(0)=1$ for co-located and co-aligned detectors.
@@
$\sigma_{IJ}^{2}(f)\approx\frac{1}{2T\Delta f}\frac{P_{I}(f)P_{J}(f)}{\gamma_{IJ}^{2}(f)S_{0}^{2}(f)},$ (15)
@@
$\beta$ Uniform[1,3][-1,3] $0.42^{+1.84}_{-0.63}$
$\log_{10}z_{s}$ Uniform[0,25][0,25] $11.3^{+8.0}_{-6.6}$
$\log_{10}z_{d}$ Uniform[0,20][0,20] $6.7^{+7.8}_{-4.9}$
$\log_{10}z_{\sigma}$ Uniform[0,18][0,18] $3.7^{+7.2}_{-3.3}$
@@
We perform a Bayesian analysis to search for the SGWB signal originating from the PBB model, using the publicly available, model-independent cross-correlation spectra $\hat{C}^{IJ}(f)$ data Abbott et al. (2021b) from the first three observing runs of Advanced LIGO and Advanced Virgo detectors. To estimate the parameters of the SGWB model arising from the PBB model, we construct a likelihood function by combining the cross-correlation spectra from all detector pairs $IJ$ Mandic et al. (2012):
[…]
where $\bm{\theta}$ represents the set of parameters characterizing the SGWB model, denoted by $\Omega_{\mathrm{M}}(f|\bm{\theta})$. The likelihood assumes that the cross-correlation spectra $\hat{C}^{IJ}(f_{k})$ follow a Gaussian distribution in the absence of a signal. The sum runs over all frequency bins $k$ and detector pairs $IJ$, with $\sigma_{IJ}^{2}(f_{k})$ being the variance of the cross-correlation statistic at each frequency bin. Using Bayes’ theorem, we express the posterior distribution of the model parameters as
$p(\bm{\theta}|C_{k}^{IJ})\propto p(C_{k}^{IJ}|\bm{\theta})\,p(\bm{\theta}),$ (17)
where $p(\bm{\theta})$ represents the prior distribution on the parameters.
[…]
[width=0.5]posts.pdf
@@
$\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}$ $=$ $\frac{p(\hat{C}^{IJ}|\text{Model with SGWB signal})}{p(\hat{C}^{IJ}|\text{Pure noise model})}$ (18)
$=$ $\frac{\int p(\hat{C}^{IJ}|\bm{\theta}_{\mathrm{GW}})\,p(\bm{\theta}_{\mathrm{GW}})\,\mathrm{d}\bm{\theta}_{\mathrm{GW}}}{\mathcal{N}}.$
The numerator denotes the marginal likelihood of the model incorporating the SGWB signal, which is calculated by integrating the likelihood $p(\hat{C}^{IJ}|\bm{\theta}_{\mathrm{GW}})$ multiplied by the prior $p(\bm{\theta}_{\mathrm{GW}})$ across the parameter space $\bm{\theta}_{\mathrm{GW}}$. The denominator $\mathcal{N}$ represents the evidence for the pure noise model, obtained by setting $\Omega_{\mathrm{M}}(f)=0$ in Eq. (16). It represents the probability of observing the data given the model that assumes only the presence of noise. The Bayes factor provides a quantitative measure of the relative support for the model with the SGWB signal compared to the pure noise model. A value of $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}>1$ indicates that the data favor the model with the SGWB signal over the pure noise model. The strength of the evidence can be interpreted using a standard scale, such as the Jeffreys scale Jeffreys (1939), where Bayes factors $\mathcal{B}_{\mathrm{NOISE}}^{\mathrm{GW}}$ exceeding 3, 10, 30, and 100 indicate substantial, strong, very strong, and decisive evidence in favor of the model with the SGWB signal, respectively.
The free parameters in our analysis are $\bm{\theta}_{\mathrm{GW}}\equiv(\beta,z_{s},z_{d},z_{\sigma})$. We list the priors of the four free parameters in Table 1. These parameters are subject to the following theoretical constraints Ben-Dayan et al. (2024):
$1\lesssim z_{\sigma}<z_{d}<z_{s},$ (19)
$\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)+\frac{3}{2}\log_{10}z_{d}-\frac{7}{2}\log_{10}z_{\sigma}\leqslant 0,$ (20)
$\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)-3\log_{10}z_{d}+\log_{10}z_{\sigma}>-42-\log_{10}4,$ (21)
$\log_{10}z_{s}<26-\log_{10}9+\frac{1}{2}\log_{10}\left(\frac{H_{1}}{M_{\mathrm{Pl}}}\right)$
$\qquad\qquad\qquad+\frac{1}{2}\left(\log_{10}z_{\sigma}-\log_{10}z_{d}\right).$ (22)
We conduct the Bayesian analysis using the Bilby package Ashton et al. (2019); Romero-Shaw et al. (2020), employing the dynamic nested sampling algorithm implemented in Dynesty Speagle (2020) with $1024$ live points to ensure adequate sampling of the parameter space.
IV Results and discussions
@@
Our analysis reveals no statistically significant evidence for the presence of a PBB SGWB signal in the Advanced LIGO and Advanced Virgo data. The Bayes factor between the PBB model and the noise-only model is found to be $0.03$, indicating “very strong” preference for the noise-only model. Consequently, we establish a lower limit for the parameter $\beta$ as $\beta\gtrsim-0.19$ at 90% confidence level. The theoretically viable range for this parameter is $0\leq\beta<3$, where the lower bound arises from the requirement of growing string coupling necessary for a smooth bounce transition Gasperini and Veneziano (2023); Conzinu et al. (2023); Modesto and Calcagni (2024), and the upper bound prevents background instabilities Kawai et al. (1998). Notably, our observational lower limit of $\beta\gtrsim-0.19$ is compatible with the theoretical constraint $\beta\geq 0$, indicating that while we do not detect a SGWB signal, our results remain consistent with the basic theoretical framework of PBB cosmology.
Our constraint $\beta\gtrsim-0.19$ at 95% C.L., while less stringent than the theoretical requirement $\beta\gtrsim 0$, provides crucial independent validation of the PBB framework. This experimental bound serves multiple important purposes: First, it offers a model-agnostic test of theoretical predictions based on string theory and stability requirements. Second, the consistency between our experimental constraint and theoretical expectations supports the validity of the underlying PBB framework—had we found $\beta\ll-0.19$, this would have challenged the entire theoretical structure. Third, our constraint from the Hz-kHz band complements and strengthens the tension revealed by PTA analyses Tan et al. (2025b); Conzinu et al. (2025), which require $\beta\approx-0.12$ to fit NANOGrav data, thereby violating theoretical bounds. This multi-band consistency demonstrates that strongly negative $\beta$ values are disfavored across different frequency ranges.
Our findings underscore the potential of GW experiments in probing alternative cosmological models and exploring the pre-big-bang era. As the sensitivity of GW detectors continues to improve, we anticipate placing even more stringent constraints on PBB models or potentially detecting the PBB SGWB signal. The upcoming observing runs of Advanced LIGO and Advanced Virgo, along with future detectors such as the Einstein Telescope Punturo et al. (2010) and Cosmic Explorer Reitze et al. (2019), will play pivotal roles in this endeavor, opening new avenues for investigating the earliest stages of the Universe and testing the fundamental principles of string theory and quantum gravity.
Acknowledgements.
QT is supported by the National Natural Science Foundation of China (Grants No. 12405055 and No. 12347111), the China Postdoctoral Science Foundation (Grant No. 2023M741148), the Postdoctoral Fellowship Program of CPSF (Grant No. GZC20240458). ZCC is supported by the National Natural Science Foundation of China under Grant No. 12405056, the Natural Science Foundation of Hunan Province under Grant No. 2025JJ40006, and the Innovative Research Group of Hunan Province under Grant No. 2024JJ1006. YW is supported by the National Natural Science Foundation of China under Grant No. 12405057. LL is supported by the National Natural Science Foundation of China (Grant No. 12505054 and  12433001).